1.Write a program, to print the following on the screen:
What percent are you aiming for in this Python module?
and then read in an answer string on the next line. It must then calculate what exam percent is required to reach this target if the user gets 90% for the Python assignments. The final mark target is calculated as 0.7*examPercent + 0.3*assignmentPercent (see Course Outline). Your last output line must be:
If you get 90% for the assignments then you will need <whatever>% for the exam



2.Write a program, that extends question 2.py to handle the requirement that the exam percent in a module must be a minimum of 40%, otherwise the module is failed.
1
Expert's answer
2015-03-18T14:09:13-0400
'''Written for Python 2.7.2.'''
print 'What percent are you aiming for in this Python module?'
while True:
try:
target = int(raw_input(''))
except:
continue
if (target >= 0) and (target <= 100):
break
else:
print 'Enter correct percentage!'
continue
exam = int((target - 0.3*90)/0.7)
print 'If you get 90% for the assignments then you will need {0}% for the exam'.format(exam)


===================================================================================
'''Written for Python 2.7.2.'''
print 'What percent are you aiming for in this Python module?'
while True:
try:
target = int(raw_input(''))
except:
continue
if (target >= 0) and (target <= 100):
exam = int((target - 0.3*90)/0.7)
if exam <40:
print 'You will receive only {0}% for exam. Try harder otherwise you will fail the module.'.format(exam)
continue
else:
print 'If you get 90% for the assignments then you will need {0}% for the exam'.format(exam)
break
else:
print 'Enter correct percentage!'
continue